# CoLa - Learning to Interactively Collaborate with Large LMs

### v2

@@ 1 Introduction
The remarkable ability of Large Language Models (LLMs) to tackle a wide range of text analysis and generation tasks without dedicated task-specific training data (Brown et al., 2020; Raffel et al., 2020; Zhang et al., 2022; Kojima et al., 2022; Yu et al., 2022; Wei et al., 2022; Wang et al., 2022) had a paradigm-changing impact on NLP. Their ability to adapt to new tasks on the fly, react to human feedback, and provide explanations for their decisions has opened the door to more natural interactions between human users and AI systems, in which the two parties can collaborate on complex multi-step tasks, such as supporting social dialogues, collaborative writing, puzzle solving (Lee et al., 2022a; b), data annotation (Kim et al., 2024), and concepts discovery (Viswanathan et al., 2024). Recent work in both the NLP and HCI communities explored these settings, investigating how to guide and evaluate the collaboration (Amershi et al., 2019; Lee et al., 2022b; Sharma et al., 2023; Collins et al., 2024b).
@@ 2 Related Work
Some of the works address the first challenge by decomposing the problem into multiple steps followed by solving each step iteratively to derive the solution  (Zhou et al., 2023; Dua et al., 2022; Wang et al., 2023; Yao et al., 2023a; b), mostly using a single LLM as a module to decompose and solve. Other works use multiple modules to separate the decomposer and solver  (Yin et al., 2024; Xu et al., 2023; Creswell et al., 2022). However, an error during any of these steps could lead into cascading errors. Our work CoLa, provides the flexibility of problem decomposition through interaction where the Guide helps by simplifying the problem for the Reasoner and our training paradigm helps to adapt the Guide based on the Reasoner’s problem solving capability and even refine or correct intermediate mistakes through interaction.
@@
Human Interaction
We draw our inspiration from Human Machine collaboration  (Collins et al., 2024a; Lee et al., 2022a; b; Collins et al., 2023; Ippolito et al., 2022; Fan et al., 2019). Some of these works show the performance of LLMs degrades on human interaction. Additionally, human interactions can be expensive. In our CoLa models we effectively simulate the role of humans, while also out-performing competitive approaches.
@@ 3.1 Interacting
We sample a handful number of examples for each domain ($<10$) and ask a human to interact with GPT-4111We use OpenAI GPT Playground to solve each problem. We share the annotation instructions and the complete process of sampling the selected examples in Appendix A.1. The human collaborator is unaware of the true solution of these problems, making this process as a true reflection of our objective. This interaction step results into forming a small seed set of conversations.
@@ 3.3 Learning
3.3.1 SFT Design
We start with a pre-trained language model Llama-3B  (Meta AI, 2024) as a starting point of the Guide model. We process the generated conversations (Sec.  3.2) and filter into successful and unsuccessful categories. A conversation is successful if it leads to the correct solution, more specifically $r_{t}$ matches to the ground truth label of the example or satisfies all the constraints in the case of the constrained text generation task.
@@ 3.3.2 RL Design
Similar to the SFT stage, we process the generated conversations (Sec. 3.2) to prepare the training set for the Critique model in the following way. We unfold all the conversations (successful and unsuccessful) into tuples of $t$ training examples, $\{(g_{1};s_{1}),(g_{1},r_{1},g_{2};s_{2}),...,(g_{1},r_{1},g_{2},...,r_{t},g_{t};s_{t})\}$, where $s_{i}$ is the label representing successful or unsuccessful in natural language text. If a conversation is successful then each $s_{i}$ is successful and vice versa. Similar to Sec. 3.3.1 we fine-tune the Critique to train on these labels.
@@
3.3.3 Ensemble Design
We further explore improving the Critique model to improve the reward signal to continue improving the interaction process. We apply the technique of ensemble for generating multiple beams of interactions between the Guide and Reasoner, such that at each interaction stage the model generates up to $K$ dialogue responses. For a given input this results into a top-$K$ beam of interactive conversations between the models. For a $k^{\text{th}}$ beam of conversation, we compute the reward for each dialogue generated by the Guide and perform the PPO update step (Sec. 3.3.2).
We also update the Critique model by fine-tuning on the generated interactions. In order to construct the training set labels for Critique we use weighted ensemble to compute the input’s label as a proxy for the ground truth label.
@@ 3.3.3 Ensemble Design
We extract the input label for a conversation, $y^{k}$ from $r_{t}^{k}$ (Reasoner’s last response) where $y$ maps to the label choice of the task, eg, $y$ would be an option for the question answering task.
Now we estimate the final label using $\operatorname*{arg\,max}$ over weighted voting of each beam, $y$:
@@
Now that we have the estimate of the input label we compare each beam’s label, $y^{k}$ against $y$ to determine if the $k^{\text{th}}$ conversation is successful or unsuccessful. This allows us to construct the training examples for the Critique as Sec. 3.3.2 ($s_{i}^{k}$) and update the parameters by performing SFT.
The rationale for the ensemble is that by generating a beam of conversations, Guide can receive more interaction episodes for the RL step and this will produce more training data for the Critique model, as well as learn to prefer multiple successful strategies, to provide better reward signal to the Guide model, such that both the models could improve through this collaboration process.
@@ 4.1 Setup
We compare CoLa against chain-of-thought style prompting approaches such as zero-CoT  (Kojima et al., 2022) and CoT  (Wei et al., 2022) for the QA domains, and instruction prompting (instruct)  (Ouyang et al., 2022) for Connections and CommonGen tasks.
@@
We benchmark two pre-trained language models - GPT-4-o-mini (GPT)  (OpenAI, 2024) and Llama 3.2 3B (Llama)  (Meta AI, 2024) as our choice for the Reasoner models. Similarly we choose GPT and Llama 3.2 3B (Llama) as our Guide models, where learning is performed for the Llama Guide model. We use Llama 3.2 1B as our Critique model. We only evaluate CoLa-RL ens for the QA domains, as the other domains have a larger space of label choices, leaving it as a future work.
@@
We implemented all the learning algorithms using TRL  (von Werra et al., 2020) in Huggingface  (Wolf et al., 2020). We use temperature  (Holtzman et al., 2020) for Guide as $0.2$ and for Reasoner as $0.1$ for all our generations.
@@
For the QA and Connections tasks, we evaluate the LLM’s response against the ground truth label of the input examples. For each system we prompt the underlying Reasoning LLM to generate the final answer based on it’s reasoning or final reasoning step (for CoLa, MAD, and refine), similar to  (Kojima et al., 2022). For the QA tasks, we compare the final answer against the ground truth option choice and compute the average percent of accuracy across the test split (Acc, ideal value: 100). For Connections the final response contains groupings of the puzzle words, and we use the following 3 metrics:
@@ 4.2 Results
We also notice that trained Guide (CoLa-SFT) boosts performance for both a stronger Reasoner (GPT) as well as smaller Reasoner (Llama), interestingly benefiting smaller Reasoner by a larger margin. We suspect this could be due to the reason that baseline performance on these QA tasks is already high, therefore less room for improvement.
@@
We observe that satisfying constraint is quite easy (Coverage), however, the challenge is to generate a more creative and coherent sentence when compared against human reference. We notice that CoLa-Prompt out-performs all the baselines except refine. Post training CoLa-SFT out-performs refine model, the difference of improvement further increases after RL phase CoLa-RL. We notice that stronger Reasoner model GPT benefits more through collaboration, which we suspect is due to the fact that output of smaller LMs is more challenging to improve  (Yu et al., 2024).
@@
Model Acc G-Acc Purity
GPT (instruct) 11.2 1.1 67.8
GPT (refine) 13.8 1.3 68.7
GPT (MAD) 15.1 1.4 70.1
GPT $\otimes$ GPT
CoLa-Prompt
15.7 1.4 71.4
Llama $\otimes$ GPT
CoLa-Prompt
10.8 1.0 62.9
Llama $\otimes$ GPT
CoLa-SFT
27.0 2.5 87.6
Llama $\otimes$ GPT
CoLa-RL
32.4 2.7 89.8
Llama (instruct) 4.7 0.6 49.3
Llama (refine) 5.8 0.7 54.6
Llama (MAD) 8.7 0.9 61.7
Llama $\otimes$ Llama
CoLa-Prompt
7.1 0.8 59.8
Llama $\otimes$ Llama
CoLa-SFT
18.1 1.8 72.8
Llama $\otimes$ Llama
CoLa-RL
20.2 1.9 74.1
Llama (sup) 11.0 1.0 64.1
Table 2: Evaluating on NYT Connections. CoLa framework for Guide $\otimes$ Reasoner.
Model Win+Tie Coverage
GPT (instruct) 51.7 98.3
GPT (refine) 59.2 99.3
GPT (MAD) 53.7 98.6
GPT $\otimes$ GPT
CoLa-Prompt
54.1 98.7
Llama $\otimes$ GPT
CoLa-Prompt
46.4 98.1
Llama $\otimes$ GPT
CoLa-SFT
60.1 99.8
Llama $\otimes$ GPT
CoLa-RL
64.7 99.9
Llama (instruct) 22.4 90.6
Llama (refine) 27.1 95.8
Llama (MAD) 24.8 92.3
Llama $\otimes$ Llama
CoLa-Prompt
24.3 92.7
Llama $\otimes$ Llama
CoLa-SFT
29.6 97.2
Llama $\otimes$ Llama
CoLa-RL
32.4 98.9
Llama (sup) 28.0 96.1
Table 3: Evaluating on CommonGen. CoLa framework for Guide $\otimes$ Reasoner.
@@ 4.3 Critique Analysis
Given the impressive performance of RL trained Guides CoLa-RL and CoLa-RL ens we perform an analysis of the role of Critique. We discuss the experiment details in App. A.3, showing the reults in Tab. 5. We observe that the Critique when fine-tuned in CoLa-RL ens, develops better ability to predict success of a strategy.
@@ 5 Human Interaction
For the QA domains, we perform an analysis by studying human interactions with a Reasoner model, Llama-3B. We compare how humans act as guides against our CoLa guides. We compare the human interactions against CoLa-Prompt and CoLa-RL ens. We provide the experiment details with results and analysis in A.4. Quantitatively we show that CoLa-RL ens Guide out-performs humans when acting as guides with the Reasoner Tab. 6. We also analyze the strategies employed by humans and CoLa models in Fig. 3 and Tab. 7. We observe that humans employ a wide range of strategies following a long tail distribution, with most of the common ones present across CoLa models as well. A key observation is that a common strategy used by humans while guiding Reasoner for QA is first providing some information and then asking the original question based on this new information (asking the original question based on new context), is completely absent in CoLa models. We believe this as a unique cognitive level understanding of the needs of a problem and assessing when enough information is available to answer the question; proposing this a motivation for future work to improve collaboration in automated systems.
Furthermore, we also observe that another common strategy shared by all the Guides is correcting intermediate reasoning steps through the interactions (correcting step in a reasoning chain). We show that CoLa-Prompt is ineffective by over-generalizing this strategy, whereas CoLa-RL ens effectively employs this strategy, even better than humans.
@@ 6 Conclusion
We propose CoLa, a novel self-supervision learning paradigm to interactively collaborate with LLMs, on a wide variety of tasks. CoLa learns to simulate humans acting as guides while collaborating with LLMs for problem solving. We train a reasonable size language model Llama-3B as the Guide and demonstrate that it can out-perform a power language model like GPT-4. We compare our approach against competitive models, perform analysis, and human study to show the effectiveness of CoLa able to strategically guide LLMs.
@@ A.2 Prompts
A.3 Role of Critique
For the QA domains, we obtain the generated response of Llama (CoT), Lllama $\otimes$ Llama CoLa-Prompt, and Lllama $\otimes$ Llama CoLa-RL ens, on the test set. In the case of CoLa the entire conversation is the generated response, whereas for Llama (CoT) the chain of though rationale with the final answer choice is the generated response. We then apply a Critique model to judge if the generated response is correct. We test three types of Critique model - critique as a pre-trained Llama-1B, and our trained Critique models obtained post SFT (SFT critique) and post RL ensemble (RL ens critique). So for our CoLa trained Critique models, the Critique response will rate the success of a conversation as a measure for correctness. Whereas off-the-shelf critique will judge the generated response for correctness.
Second as the Critique is fine-tuned on CoLa conversations, it develops a better ability to judge the responses against the ground truth, with performance increasing for CoLa-RL ens model. This demonstrates that fine-tuning leads to better responses further leads to better critique ability, hence improving both the Guide and Critique models.
Model + Critique SocialIQA CSQA2
P R F1 P R F1
Llama + critique 0.91 0.85 0.88 0.86 0.79 0.82
Table 5: Critique Success. CoLa - Guide $\otimes$ Reasoner
A.4 Human Analysis
We sample a few examples from the test split of the QA tasks, around 200 examples from CSQA2 and 50 examples from SocialIQA, and ask humans to interact with LLama-3B in a similar style as CoLa. We share the same instructions and guidelines as mentioned in App A.1 and analyze their responses. Tab.  6 shows the performance of different interactive systems when using different Guide with the same Reasoner model. We also compute the average number of interaction rounds (Rounds), and relevance (Rel) of each interaction by prompting GPT-4 with the contextualized conversation and ground truth label to ask if the given Guide interaction is relevant or not. We observe that CoLa-SFT and CoLa-RL ens out-perform humans when acting as Guide, a positive result of learning to interact effectively.
Figure 3: Strategy Pie Chart

### v3

# CoLa: Learning to Interactively Collaborate with Large Language Models

@@ 1 Introduction
The remarkable ability of Large Language Models (LLMs) to tackle a wide range of text analysis and generation tasks without dedicated task-specific training data (Brown et al., 2020; Raffel et al., 2020; Zhang et al., 2022; Kojima et al., 2022; Yu et al., 2022; Wei et al., 2022; Wang et al., 2022) had a paradigm-changing impact on NLP. Their ability to adapt to new tasks on the fly, react to human feedback, and provide explanations for their decisions has opened the door to more natural interactions between human users and AI systems, in which the two parties can collaborate on complex multi-step tasks, such as supporting social dialogues, collaborative writing, puzzle solving (Lee et al., 2022a; b), data annotation (Kim et al., 2024), and concepts discovery (Viswanathan et al., 2024; Pujari & Goldwasser, 2025). Recent work in both the NLP and HCI communities explored these settings, investigating how to guide and evaluate the collaboration (Amershi et al., 2019; Lee et al., 2022b; Pacheco et al., 2023; Sharma et al., 2023; Collins et al., 2024b).
@@ 2 Related Work
Human Interaction
We draw our inspiration from Human Machine collaboration  (Collins et al., 2024a; Lee et al., 2022a; b; Collins et al., 2023; Ippolito et al., 2022; Fan et al., 2019), such that humans interact with machine (language models in our settings), with the objective of collaboration for solving a task. However, the key challenge is effectively scaling human interactions, as obtaining human input is expensive. Furthermore,  Lee et al. (2022b) notes that using human interaction can also degrade the task performance. We design CoLa by building on only handful human interaction examples and effectively scaling them by designing training objectives which model successful collaboration, on a wide range of tasks.
@@
Some of the works address the first challenge by decomposing the problem into multiple steps followed by solving each step iteratively to derive the solution  (Zhou et al., 2023; Dua et al., 2022; Wang et al., 2023; Yao et al., 2023a; b), mostly using a single LLM as a module to decompose and solve. Other works use multiple modules to separate the decomposer and solver  (Yin et al., 2024; Xu et al., 2023; Creswell et al., 2022). However, an error during any of these steps could lead into cascading errors. Our work CoLa, provides the flexibility of problem decomposition through interaction where the Guide helps by simplifying the problem for the Reasoner and our training paradigm helps to adapt the Guide based on the Reasoner’s problem solving capability and even refine or correct intermediate mistakes through interaction. Furthermore, CoLa supports interactive collaboration between agents – Guide and Reasoner through dialogue acts not plans, drawing this design inspiration from human-machine collaboration.
@@ 3.1 Interacting
We sample a handful number of examples for each domain (<10<10) and ask a human to interact with GPT-4111We use OpenAI GPT Playground to solve each problem. We share the annotation instructions and the complete process of sampling the selected examples in Appendix A.1. The human collaborator is unaware of the true solution of these problems, making this process as a true reflection of our objective. This interaction step results into forming a small seed set of conversations.
@@ 3.3 Learning
SFT Design
We start with a pre-trained language model Llama-3B  (Meta AI, 2024) as a starting point of the Guide model. We process the generated conversations and filter into successful and unsuccessful categories. A conversation is successful if it leads to the correct solution, more specifically $r_{t}$ matches to the ground truth label of the example or satisfies all the constraints in the case of the constrained text generation task.
@@
Similar to the SFT stage, we process the generated conversations to prepare the training set for the Critique model in the following way. We unfold all the conversations (successful and unsuccessful) into tuples of $t$ training examples, $\{(g_{1};s_{1}),(g_{1},r_{1},g_{2};s_{2}),...,(g_{1},r_{1},g_{2},...,r_{t},g_{t};s_{t})\}$, where $s_{i}$ is the label representing successful or unsuccessful in natural language text. If a conversation is successful then each $s_{i}$ is successful and vice versa. Similar to SFT we fine-tune the Critique to train on these labels.
@@
Ensemble Design
We further explore improving the Critique model to continue improving the interaction process. We design this iterative improvement loop using the technique of ensemble. We draw the inspiration from the EM (expectation-maximization) algorithm, such that in the E-step we generate multiple beams of conversations between the Guide and Reasoner, and score each conversation based on the reward generated by the Critique model, to compute the expected posterior label using the reward score as the weighted ensemble. Whereas, in the M-step we use the previously generated conversations to update the parameters of the Critique model, based on the reward score of the conversations.
During the start of the conversation we prompt the Reasoner to generate top-$K$ responses. Each response is then provided to the Guide and the interactions are run in-parallel, creating $K$ conversations thereafter. For a $k^{\text{th}}$ beam of conversation, we compute the reward for each dialogue generated by the Guide and perform the PPO update step.
We also update the Critique model by fine-tuning on the generated interactions. In order to construct the training set labels for Critique we use weighted ensemble to estimate (E-step) the input’s label as a proxy for the ground truth label.
@@
We extract the input label for a conversation, $y^{k}$ from $r_{t}^{k}$ (Reasoner’s last response) where $y$ maps to the label choice of the task, eg, $y$ would be an option for the question answering task. Now we estimate the final label using $\operatorname*{arg\,max}$ over weighted voting of each beam, $y$:
@@
Now that we have the estimate of the input label we compare each beam’s label, $y^{k}$ against $y$ to determine if the $k^{\text{th}}$ conversation is successful or unsuccessful. This allows us to construct the training examples for the Critique ($s_{i}^{k}$) and update the parameters by performing SFT (M-step). The rationale is that by generating $K$-fold conversations for each problem, results in multiple rationales as solutions to the original problem, providing a rich training data for the Critique model. This yields a better reward signal for the Guide model for the RL step, such that both the models could improve through this iterative process.
@@ 4.1 Setup
We compare CoLa against chain-of-thought style prompting approaches such as zero-CoT  (Kojima et al., 2022) and CoT  (Wei et al., 2022) for the QA domains, and instruction prompting (instruct)  (Ouyang et al., 2022) for Connections and CommonGen.
@@
We benchmark two pre-trained language models - GPT-4-o-mini (GPT)  (OpenAI, 2024) and Llama 3.2 3B instruct-chat (Llama)  (Meta AI, 2024) as our choice for the Reasoner models. Similarly we choose GPT and Llama 3.2 3B instruct-chat (Llama) as our Guide models, where learning is performed for the Llama Guide model. We use Llama 3.2 1B instruct-chat as our Critique model. We only evaluate CoLa-RL ens for the QA domains, as the other domains have a larger space of label choices, leaving it as a future work.
@@
We implemented all the learning algorithms using TRL  (von Werra et al., 2020) in Huggingface  (Wolf et al., 2020). We use temperature  (Holtzman et al., 2020) for Guide as $0.2$ and for Reasoner as $0.1$ for all our generations. For CoLa-RL ens, we use $K=5$ and $K=3$ as the top-$K$ for SocialIQA and CSQA2 respectively.
@@
For the QA and Connections tasks, we evaluate the LLM’s response against the ground truth label of the input examples. For each system we prompt the underlying Reasoning LLM to generate the final answer based on it’s reasoning or final reasoning step (for CoLa, MAD, and refine), similar to  (Kojima et al., 2022). For the QA tasks, we compare the final answer against the ground truth option choice and compute the average percent of accuracy across the test split (Acc, ideal value: 100).
For Connections the final response contains groupings of the puzzle words, and we use the following 3 metrics:
@@ 4.2 Results
We also notice that a trained Guide (CoLa-SFT) boosts performance for both a stronger Reasoner (GPT) as well as smaller Reasoner (Llama), interestingly benefiting smaller Reasoner by a larger margin. We suspect this could be due to the reason that baseline performance on these QA tasks is already high, therefore less room for improvement.
@@
We show the results for the constrained text generation task in Tab. 2. We observe that satisfying constraint is quite easy (Coverage), however, the challenge is to generate a more creative and coherent sentence when compared against human reference. We notice that CoLa-Prompt out-performs all the baselines except refine. Post training CoLa-SFT out-performs the refine model, the difference of improvement further increases after RL phase CoLa-RL. We notice that stronger Reasoner model GPT benefits more through collaboration, which we suspect is due to the fact that output of smaller LMs is more challenging to improve  (Yu et al., 2024).
@@
Model Connections CommonGen
Acc G-Acc Purity Win+Tie Coverage
GPT (instruct) 11.2 1.1 67.8 51.7 98.3
GPT (refine) 13.8 1.3 68.7 59.2 99.3
GPT (MAD) 15.1 1.4 70.1 53.7 98.6
GPT $\otimes$ GPT CoLa-Prompt 15.7 1.4 71.4 54.1 98.7
Llama $\otimes$ GPT CoLa-Prompt 10.8 1.0 62.9 46.4 98.1
Llama $\otimes$ GPT CoLa-SFT 27.0 2.5 87.6 60.1 99.8
Llama $\otimes$ GPT CoLa-RL 32.4 2.7 89.8 64.7 99.9
Llama (instruct) 4.7 0.6 49.3 22.4 90.6
Llama (refine) 5.8 0.7 54.6 27.1 95.8
Llama (MAD) 8.7 0.9 61.7 24.8 92.3
Llama $\otimes$ Llama CoLa-Prompt 7.1 0.8 59.8 24.3 92.7
Llama $\otimes$ Llama CoLa-SFT 18.1 1.8 72.8 29.6 97.2
Llama $\otimes$ Llama CoLa-RL 20.2 1.9 74.1 32.4 98.9
Llama (sup) 11.0 1.0 64.1 28.0 96.1
Table 2: Evaluation on NYT Connections (Acc, G-Acc, Purity) and CommonGen (Win+Tie, Coverage) using CoLa framework for Guide $\otimes$ Reasoner.
@@ 4.3 Critique Analysis
Given the impressive performance of the CoLa models we perform an analysis of the role of Critique. For the QA domains, we obtain the generated response of Llama (CoT), Llama $\otimes$ Llama CoLa-Prompt, and Llama $\otimes$ Llama CoLa-RL ens, on the test set. We then apply a Critique model to judge if the generated response is correct. We test three types of Critique model - critique as a pre-trained Llama-1B, and our trained Critique models obtained post SFT (SFT critique) and post RL ensemble (RL ens critique).
Second as the Critique is fine-tuned on CoLa conversations, it develops a better ability to judge the responses against the ground truth, with performance increasing for CoLa-RL ens model.
Model + Critique SocialIQA CSQA2
P R F1 P R F1
Llama + critique 0.91 0.85 0.88 0.86 0.79 0.82
Table 3: Critique Success. CoLa - Guide $\otimes$ Reasoner
@@ 5 Human Interaction
We conduct a study to understand how effective is our CoLa simulated Guide when compared to actual human guides. Our primary goal is to observe the similarities and differences between the interactions of a human vs a simulated human with an LLM. For this we propose an interactive experiment between Guide and Reasoner. We fix the Reasoner model to Llama-3B, and the Guide could either be a human or CoLa Guide. We provide the experiment details in  A.3. We analyze the interactions of humans with the Reasoner model and compare against the CoLa interactions.
Human vs. CoLa: who is a better guide? We observe the quantitative results in Tab. 5. The results positively demonstrate how training a Guide is more effective at guidance when compared against humans. However, these QA tasks itself are not challenging for humans, so we further investigate the interactions to discover the reason for humans not out-performing the CoLa guides. We observe that humans mostly have an idea of the correct answer, however, they interact using complex reasoning strategies on which the Reasoner finds challenging to adapt. Whereas, the CoLa guides were specifically trained with the Reasoner model, making the adaptation process easier.
Do Humans and CoLa guide differently? Our next experiment compares the fine grain strategies employed by humans vs the CoLa guides. We select the CSQA2 task and qualitatively analyze the strategies. We sample few interactions and come up with a strategy for each sampled interaction. We then use these as few-shot exemplars to generate the remaining strategies for each interaction, by prompting GPT4-o-mini (OpenAI, 2024).
We show the distribution of these strategies for each Guide in Fig 3 and Tab 6 – with the frequency of a strategy along with its correctness i.e. how often does a given interaction’s strategy lead to the correct answer at the end of the conversation. We observe that humans employ a wide range of strategies following a long tail distribution, with most of the common ones present across CoLa models as well. A common strategy shared by all the Guides is correcting intermediate reasoning steps through the interactions (correcting step in a reasoning chain). We show that CoLa-Prompt is ineffective by over-generalizing this strategy, whereas CoLa-RL ens effectively employs this strategy, even better than humans. A key observation is that a common strategy used by humans while guiding Reasoner for QA is first providing some information and then asking the original question based on this new information (asking the original question based on new context), is completely absent in CoLa models. We believe this as a unique cognitive level understanding of the needs of a problem and assessing when enough information is available to answer the question; proposing this a motivation for future work to improve collaboration in automated systems.
Figure 3: CSQA2 Interactions Strategy Pie Chart
@@ 6 Conclusion
Drawing inspiration from human machine collaboration we propose CoLa, a framework to simulate human guidance for LLMs with the goal of solving a wide range of language tasks. CoLa provides guidance to LLMs by generating dialogues and engaging in a multi-turn conversation leading to the task solution. We discuss various training approaches to learn to simulate the human guidance, and propose a novel self-supervision learning paradigm based on Reinforcement Learning, which out-performs competitive baselines. We also conduct a human analysis to study the differences between human guidance and CoLa guidance, observing that humans employ wide range of guidance strategies, whereas CoLa when trained can out-perform human guidance and perform some of the guidance strategies better than humans.
Acknowledgments
We are very grateful to all the members of the Purdue NLP lab, for their help with the annotations and valuable suggestions for CoLa. We also thank the reviewers for their feedback for improving the paper. This paper was partially supported by an NSF CAREER award IIS-2048001.
@@ A.2 Prompts
A.3 Human Analysis
We sample a few examples from the test split of the QA tasks, around 200 examples from CSQA2 and 50 examples from SocialIQA, and ask humans to interact with LLama-3B in a similar style as CoLa. We share the same instructions and guidelines as mentioned in App A.1 and analyze their responses. Tab. 5 shows the performance of different interactive systems when using different Guide with the same Reasoner model. We also compute the average number of interaction rounds (Rounds), and relevance (Rel) of each interaction by prompting GPT-4 with the contextualized conversation and ground truth label to ask if the given Guide interaction is relevant or not. We observe that CoLa-SFT and CoLa-RL ens out-perform humans when acting as Guide, a positive result of learning to interact effectively.